OBSERVATION OF ORTHO-PARA H_{3}^{+} SELECTION RULES IN PLASMA CHEMISTRY. A TIME DEPEN-DENT STUDY IN A HOLLOW CATHODE DISCHARGE.

M. CORDONNIER, Laboratoire de Spectroscopie Hertzienne, Associé au C.N.R.S, U.F.R de Physique, Université des Sciences et Technologies de Lille, Bât. P5, Villeneuve d’Ascq, 59655, France; D. UY, JILA, University of Colorado at Boulder, Boulder, CO, 80309; Y. ZHANG and T. OKA, Department of Chemistry and Department of Astronomy and Astrophysics, The University of Chicago, Chicago, IL, 60637.

During the past two Columbus meetings, we reported our observation of a nonthermal ortho-to-para distribution of H_{3}^{+} in hydrogen plasmas, where we found selection rules on nuclear spin modifications to hold in a chain of plasma chemical reactions. The solution of rate equations incorporating nuclear spin branching ratios gave an ortho-to-para ratio of H_{3}^{+} in good agreement with experiments. However, our observations were limited to steady-state plasma conditions, and no direct information relative to the kinetics of the plasma could be derived. For this reason, a new time dependent experiment has been carried out in a hydrogen plasma. Our study consisted of monitoring the evolution of ortho- and para-H_{3}^{+} absorptions during and after a short pulse discharging normal H_{2} or para-H_{2} confined in a hollow cathode. Deviations from thermal conditions were found stronger than those observed in the steady-state work, and the study of the H_{3}^{+} absorption signals allowed a more detailed analysis of the chemical processes occurring during the thermalization of the o-H_{3}^{+} and p-H_{3}^{+} in the cell. An experimental value for the ratio of rate constants of the proton hop (H̃_{2} + H_{3}^{+} → H̃_{2}H^{+} + H_{2}) and the hydrogen exchange (H̃_{2} + H_{3}^{+} → H_{2}H̃^{+} + HH̃) reactions has been obtained to explain the ortho-to-para H_{3}^{+} ratios observed in the pulsed plasma.